OpenStudy (anonymous):

Given the equation Square root of 2x plus 1 = 3, solve for x and identify if it is an extraneous solution. ^ I really don't understand this topic, whatsoever. Could someone help?

OpenStudy (anonymous):

\[\sqrt{2x+1}=3\] right?

OpenStudy (anonymous):

get rid of the radical by squaring both sides you get \[2x+2=3^2\] i.e. \[2x+1=9\] and solve that one for \(x\) in two steps

OpenStudy (anonymous):

typo there, sorry, i meant you get \(2x+1=3^2\) i.e. \(2x+1=9\)

OpenStudy (anonymous):

thank you, so the answer is 4, but what exactly does extraneous mean? two of the answer choices are 4, but one says the solution is extraneous and one says that it isn't.

OpenStudy (anonymous):

the answer is 4, and there are no extraneous solutions, that is the one and only answer

OpenStudy (anonymous):

i will give you an example of an extraneous solution if you like when you square both sides, you could introduce a solution to the squared equation that was not a solution to the original equation that did not happen in this case

OpenStudy (anonymous):

okay thanks; so if it were extraneous that's basically saying that there's more than possible one answer for x?

hartnn (hartnn):

if you plug back x=4 in your original equation, you will indeed get left side = right side. so there's no extraneous solution. if you got some value and after plugging it back into original equation, you do not get left side = right side then that solution is extraneous

OpenStudy (anonymous):

you get an extraneous solution sometimes when you square both sides and end up with a quadratic equation that has two solutions, one of them could be wrong

OpenStudy (anonymous):

here is a really really simple example the solution to \(x=-3\) is \(-3\) but if you square both sides you get \(x^2=9\) which, when solving, gives two solutions \(\{-3,3\}\) but only \(-3\) is a solution to the original equation

OpenStudy (anonymous):

i THINK i understand where you getting it.. okay; i have two of those kinds of questions left. if i give you what i think is the answer, can you like, tell me if i got it right? lol .-.

OpenStudy (anonymous):

i can give you another example of one with an extraneous solution if you like \[\sqrt{x-1}=2x-3\]

OpenStudy (anonymous):

ok go ahead btw if you have answer choices, the easiest thing to do is check the numbers and see which ones work

OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11. x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous I chose the first one? I got x equals 21, but when I plugged it back into square root of x minus 5 + 7 = 11 it was wrong? sigh .-,

hartnn (hartnn):

x=21 is correct when you plugged it back in, how did you get it wrong, try again ?

hartnn (hartnn):

what is 21- 5 ?

OpenStudy (anonymous):

16 but plus 7 is 23?

hartnn (hartnn):

but there's a square root too! \(\sqrt{x-5} = \sqrt{21-5}=\sqrt 16 = 4\)

hartnn (hartnn):

and whats 4+7 ? ;)

OpenStudy (anonymous):

oh! lol okay, so then that's not an extraneous solution?

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

okaayy, my last question? Given the equation −4 Square root of x minus 3 = 12, solve for x and identify if it is an extraneous solution. x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous not gonna lie I don't know where to start for this one since I don't see an example like this in my lesson.

OpenStudy (anonymous):

you can make me work it, just kinda like give me a start

hartnn (hartnn):

start by dividing both sides by -4

hartnn (hartnn):

then, you can square both sides

OpenStudy (anonymous):

the -4 turns into a 1, the 12 turns into -3

hartnn (hartnn):

correct

hartnn (hartnn):

now square both sides

OpenStudy (anonymous):

the 1 is gonna stay a 1 though right?

OpenStudy (anonymous):

-3 will turn into 9

hartnn (hartnn):

yess

OpenStudy (anonymous):

x = 12

hartnn (hartnn):

now check whether it is extraneous ?

OpenStudy (anonymous):

it is extraneous, right?

hartnn (hartnn):

absolutely! now you are good at these ;)

OpenStudy (anonymous):

ayeeeeeee! omg thank you so much for the help! lol including @satellite73

hartnn (hartnn):

glad i could help :) welcome ^_^